Week of 200330

WLCG Operations Call details

  • For remote participation we use the Vidyo system. Instructions can be found here.

General Information

  • The purpose of the meeting is:
    • to report significant operational issues (i.e. issues which can or did degrade experiment or site operations) which are ongoing or were resolved after the previous meeting;
    • to announce or schedule interventions at Tier-1 sites;
    • to inform about recent or upcoming changes in the experiment activities or systems having a visible impact on sites;
    • to provide important news about the middleware;
    • to communicate any other information considered interesting for WLCG operations.
  • The meeting should run from 15:00 Geneva time until 15:20, exceptionally to 15:30.
  • The SCOD rota for the next few weeks is at ScodRota
  • General information about the WLCG Service can be accessed from the Operations Portal
  • Whenever a particular topic needs to be discussed at the operations meeting requiring information from sites or experiments, it is highly recommended to announce it by email to wlcg-scod@cernSPAMNOTNOSPAMPLEASE.ch to allow the SCOD to make sure that the relevant parties have the time to collect the required information, or invite the right people at the meeting.

Best practices for scheduled downtimes

Monday

Attendance:

  • remote: Kate (chair, WLCG, DB), Maarten (WLCG, ALICE), Julia (WLCG), Borja (monitoring), Vladimir (LHCb), Michal (ATLAS), Olga (computing), Roberto (storage), Darren (RAL), Andrew (TRIUMF), Vincent (security), Onno (NL-T1), David (IN2P3), Xin (BNL), Jens (NDGF), Dave (FNAL)

Experiments round table:

  • ATLAS reports ( raw view) -
    • Activities:
      • Ongoing reprocessing
      • SFO->CTA stress test - additional testing done to debug low throughput problem
    • Issues
      • 11 RAW files were found to be corrupted on IN2P3-CC - recovered
      • looping jobs at CNAF
        • caused by few bugs in pilot
        • fixes to be released today
      • job problems related to switch to daylight saving time on Sunday
        • sites could observe a peak of lost heartbeats in the night
        • some sites could see less jobs as aCT service crashed and was restarted Sunday morning

  • CMS reports ( raw view) -
    • This week is CMS week: Due to the plenary being parallel to the WLCG call nobody from CMS will join the call
    • No big problems to report

  • ALICE -
    • Large numbers of jobs have been using excessive amounts of memory today
      • All due to particular analysis trains, possibly badly tested
      • The relevant experts have been asked to look into this matter ASAP
      • We apologize for the problems that some sites suffered due to this!

  • LHCb reports ( raw view) -
    • Activity:
      • Stripping.
      • Staging for next stripping round.
    • Issues:

Sites / Services round table:

  • ASGC: nc
  • BNL: Lab is in minimal operations mode. DCache upgrade will be delayed as it requires HW reshuffling
  • CNAF: NTR
  • EGI: nc
  • FNAL: NTR
  • IN2P3: 110 TB ALICE disk server has been down since yesterday evening. Due to current access restrictions, sysadmins will check tomorrow if Dell intervention is necessary or not.
  • JINR: NTR
  • KISTI: nc
  • KIT: Reminder for storage downtime on 1st of April, where we have to update the firmware of our storage systems. All storage will be offline, so all VOs will be affected!
  • NDGF: NTR
  • NL-T1: Surfsara dCache upcoming downtime 20 to 24 April https://goc.egi.eu/portal/index.php?Page_Type=Downtime&id=28583
  • NRC-KI: nc
  • OSG: nc
  • PIC: We still have some errors for LHCb transfers at PIC. Having a look to it. From tomorrow on, only essential services in Spain are allowed to work. We will continue operating PIC remotely.
  • RAL: NTR
  • TRIUMF: NTR

  • CERN computing services: OTG0055697 HTCondor schedds downtime due to network intervention, intervention on 1st April.
  • CERN storage services: NTR
  • CERN databases: NTR
  • GGUS: NTR
  • Monitoring:
    • On the 1st of April REBUS will be made read only, this is the first step towards the migration of REBUS into CRIC
  • MW Officer: nc
  • Networks: NTR
  • Security: NTR

AOB:

Edit | Attach | Watch | Print version | History: r19 < r18 < r17 < r16 < r15 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r19 - 2020-03-30 - KateDziedziniewicz
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    LCG All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright & 2008-2020 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding TWiki? Send feedback